# Seeing Through Deception: Uncovering Misleading Creator Intent in Multimodal News with Vision-Language Models

### v2

@@ Abstract
The real-world impact of misinformation stems from the underlying misleading narratives that creators seek to convey. As such, interpreting misleading creator intent is essential for multimodal misinformation detection (MMD) systems aimed at effective information governance. In this paper, we introduce an automated framework that simulates real-world multimodal news creation by explicitly modeling creator intent through two components: the desired influence and the execution plan. Using this framework, we construct DeceptionDecoded, a large-scale benchmark comprising 12,000 image-caption pairs aligned with trustworthy reference articles. The dataset captures both misleading and non-misleading intents and spans manipulations across visual and textual modalities. We conduct a comprehensive evaluation of 14 state-of-the-art vision-language models (VLMs) on three intent-centric tasks: (1) misleading intent detection, (2) misleading source attribution, and (3) creator desire inference. Despite recent advances, we observe that current VLMs fall short in recognizing misleading intent, often relying on spurious cues such as superficial cross-modal consistency, stylistic signals, and heuristic authenticity hints. Our findings highlight the pressing need for intent-aware modeling in MMD and open new directions for developing systems capable of deeper reasoning about multimodal misinformation.  [^1] Link to DeceptionDecoded: [GitHub Repo].
Seeing Through Deception: Uncovering Misleading Creator Intent in
Multimodal News with Vision-Language Models
Jiaying Wu1Fanxiao Li2Min-Yen Kan1Bryan Hooi1 1National University of Singapore   2Yunnan University jiayingwu@u.nus.edu, lifanxiao@mail.ynu.edu.cn, {kanmy, bhooi}@comp.nus.edu.sg
Content Warning: this paper contains
potentially harmful text and images.
@@ 1 Introduction
Figure 1: The importance of detecting misleading creation intent in multimodal news beyond surface-level cross-modal consistency. For instance, a malicious creator can craft captions that are semantically consistent with the image yet deliberately convey a false narrative.
@@
Multimodal misinformation, which combines persuasive text with compelling visuals, poses significant threats to public understanding and can lead to serious societal harm Zhou and Zafarani (2020); Alam et al. (2022); Do Nascimento et al. (2022). A growing body of research in multimodal misinformation detection (MMD) focuses on identifying cross-modal misalignment, particularly in two major forms: (1) out-of-context (OOC) misinformation Yuan et al. (2023); Qi et al. (2024), where images and captions from unrelated events or time periods are falsely paired; and (2) multimodal media manipulation  Shao et al. (2023); Liu et al. (2025), involving subtle changes such as shifts in facial expressions or caption phrasing that alter the perceived message. However, current MMD benchmarks often simulate such misalignments in artificial ways, either by mismatching image and text with CLIP-based Radford et al. (2021) similarity scores Luo et al. (2021) or through sentiment word substitutions Sudhakar et al. (2019); Shao et al. (2023). These arbitrary strategies create a fundamental gap between constructed benchmarks and the complexity of real-world MMD.
@@
Beyond surface-level cross-modal misalignment, the key to effective real-world misinformation governance lies in detecting and understanding misleading creator intent Appelman et al. (2022); Jaidka et al. (2025). Many misinformation campaigns are deliberately crafted to advance specific agendas, often without the audience’s awareness, yet still manage to substantially influence public opinion Ecker et al. (2022); Broda and Strömbäck (2024). For example, as illustrated in Figure 1, a slight modification to a news caption can lead readers to attribute significant casualties to government negligence, ultimately eroding institutional trust. Although preliminary efforts have explored intent interpretation Da et al. (2021); Gabriel et al. (2022); Wang et al. (2023), they are often limited to unimodal settings and fail to capture the synergy between visual and textual cues. Moreover, these approaches rely on reader inference of creator intent, which are prone to error due to the deceptive framing of such content. These limitations highlight a pressing need for transparent modeling of ground-truth creator intents in multimodal content, along with systematic benchmarks for detecting misleading communicative objectives.
@@
To this end, we establish an intent-guided framework for simulating intent-aware multimodal news reporting. Specifically, we define diverse combinations of desired influence and execution plans, two core components that characterize real-world communication strategies Hallahan et al. (2007); Paul (2011). Building on this, we introduce DeceptionDecoded, a large-scale benchmark designed to reveal misleading creator intents in multimodal news. Constructed upon real-world image-caption pairs from the VisualNews Liu et al. (2021) repository, DeceptionDecoded contains 12,000 instances created with misleading and non-misleading intent, spanning a broad spectrum of misleading intent types and framing strategies. Each instance is paired with a trustworthy reference article to serve as contextual grounding, and the quality of the generated data are further validated through human evaluation (Section 3.4). The benchmark supports evaluation of three intent-centric tasks related to MMD: (1) Misleading Intent Detection, which determines whether a news piece was intentionally crafted to mislead; (2) Misleading Source Attribution, which identifies whether the misleading signal originates from the image or the text; and (3) Creator Desire Inference, which infers the creator’s intended area of societal impact, such as political polarization or public health disruption.
Using DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs). Despite recent advances in multimodal reasoning Hurst et al. (2024); Anthropic (2025c), we find that state-of-the-art VLMs still struggle with intent understanding in the context of multimodal news. For instance, even relatively strong models such as GPT-4o-mini OpenAI (2024) achieve only 18.8% accuracy when misleading intent is subtly expressed through caption manipulation. More concerningly, our analysis reveals that VLMs tend to rely on spurious features when detecting misleading intent, including superficial image–text consistency, stylistic patterns in the text, and heuristic cues suggesting authenticity. This over-reliance leads to significant performance degradation under simple adversarial manipulations, such as stylistic reframing Chen and Shu (2024); Wu et al. (2024) and persuasive prompting attacks Zeng et al. (2024). Our benchmark and empirical analyses establish a foundation for future MMD research that targets deeper semantic reasoning and intent-aware understanding beyond shallow cross-modal consistency patterns.
@@
Benchmark Task Setup Content Modality Creator Intent News Context
@@ 2 Related Work
2.1 Multimodal Misinformation Detection
With the increasing prevalence of persuasive visual–textual misinformation on online platforms, research has expanded from detecting purely textual misinformation Rashkin et al. (2017); Chen and Shu (2024); Wu et al. (2024) to tackling multimodal misinformation detection (MMD). A growing body of work emphasizes misinformation arising from subtle cross-modal misalignments, such as out-of-context (OOC) misinformation Qi et al. (2024) and multimedia manipulation Shao et al. (2023); Liu et al. (2025). Recent advances leverage VLMs like LLaVA Liu et al. (2023) and GPT-4o Hurst et al. (2024) within retrieval-augmented frameworks, where the model reasons over multimodal inputs using retrieved trustworthy evidence Khaliq et al. (2024); Xuan et al. (2024); Zhou et al. (2024); Braun et al. (2024). While effective in grounding responses, these approaches largely overlook the role of news creation intent, which has been identified as a key driver of misinformation Sharma et al. (2019). Our work builds upon this evidence-based paradigm by evaluating VLMs on their ability to detect misleading communicative intent in multimodal news. We show that even the most advanced models still fall short in this setting, highlighting the need for intent-aware frameworks in MMD.
@@ 2.2 Misinformation Benchmarks
Various benchmarks has been developed to support misinformation-related research Hanselowski et al. (2019); Yao et al. (2023). Representative datasets such as PolitiFact and GossipCopShu et al. (2020) collect naturally occurring misinformation on social media. Other works approach MMD from the perspective of multimedia manipulation; for example, NewsCLIPpings Shu et al. (2020) and DGM4 Shao et al. (2023) introduce cross-modal mismatches by substituting images or captions based on visual or textual similarity. Mixed-source datasets such as Fakeddit Nakamura et al. (2020) and MMFakeBench Liu et al. (2025) include both naturally occurring and synthetically manipulated content. While preliminary efforts such as EMU Da et al. (2021), NewsINT Wang et al. (2023), and MRF Gabriel et al. (2022) have considered the role of news creator intent, they are limited to unimodal misinformation. Moreover, the notion of creator intent in these datasets is typically annotated from readers’ subjective perspective, which may misalign with the actual communicative purpose. In this work, we introduce DeceptionDecoded, a benchmark featuring creator-produced multimodal misinformation, where intent is explicitly defined during content generation. A comparison of DeceptionDecoded to prior work is shown in Table 1.
Figure 2: DeceptionDecoded: Overview of multimodal news curation guided by diverse simulated news creator intents, including both misleading and non-misleading cases.
@@ 3 DeceptionDecoded
To systematically investigate the role of misleading creator intent in multimodal news, it is crucial to construct a large-scale benchmark that explicitly defines diverse communicative intents across various news domains and uses these intents to guide content generation. However, this is challenging to achieve with real-world news data, where content is already published and annotators can only infer intent retrospectively from a reader’s perspective. To overcome this, we introduce DeceptionDecoded (overviewed in Figure 2), a dataset of 12,000 multimodal news instances generated through an intent-guided synthesis pipeline. Each instance is grounded in a predefined creator intent setup, enabling controlled formulation while preserving content realism. Data quality is validated through rigorous human evaluation.
@@ 3.1 Source News Collection
We draw from VisualNews Liu et al. (2021), a large-scale repository of trustworthy multimodal news, to ensure factual grounding. Specifically, we select 10 topic categories that offer broad societal relevance and sample richness. These span diverse domains such as politics, disasters, and public health. The full list of selected topics is provided in Appendix A.1.
To support meaningful evaluation of misinformation detection, we apply rigorous filtering criteria guided by principles from both misinformation governance and professional journalism. From a governance perspective, we prioritize content that serves the public interest Kruger et al. (2024), focusing on events with high societal relevance and potential for wide public impact. In line with journalistic standards, we require that selected content demonstrates professionalism, neutrality, and clarity Maras (2013); Farley et al. (2014), ensuring that language is objective, free from bias, and unambiguous. To further mitigate harm, we exclude instances that reference specific individuals or identifiable public figures (detailed in Appendix A.2).
This process yields 2,000 high-quality samples, evenly distributed across the 10 selected topics. Each instance is structured as $N=\{I,T,A\}$, where $I$ is the news image, $T$ is the accompanying caption, and $A$ is a trusted reference article.
@@ 3.2 Creator Intent Establishment
Given the context provided by each original news instance $N=\{I,T,A\}$, we simulate both malicious and trustworthy content creators using GPT-4o Hurst et al. (2024), leveraging its demonstrated capacity for social simulation Anthis et al. (2025).
Inspired by strategic communication theories Hallahan et al. (2007); Paul (2011) and prior work on textual intent inference Wang et al. (2023), we conceptualize creator intent ($C_{\text{int}}$) along two core dimensions: (1) desired influence and (2) execution plan. The desired influence refers to the societal sector(s) the creator aims to affect, selected from a predefined list of options (up to three per instance; see Appendix A.3). The execution plan is generated as open-ended text, detailing how the creator intends to achieve the intended influence.
@@ 3.3 Intent-Guided Multimodal News Creation
Guided by $C_{\text{int}}$, we generate misleading and non-misleading variants of each news piece obtained in Section 3.1, by either modifying the image or text. Misleading intents are categorized into two levels: (1) subtle (e.g., minor distortions to framing or background details to nudge reader interpretation) and (2) significant (e.g., major alterations that substantially change the perceived message). For non-misleading cases, we prompt the creator to faithfully paraphrase the caption or reconstruct the image while retaining alignment with the article.
For textual modification, we use GPT-4o Hurst et al. (2024) to generate captions aligned with the specified intent, as defined in Section 3.2, while keeping the original image unchanged. This produces a sample of the form $N_{\text{text}}=\{I,T^{\prime},A\}$, where $I$ is the original image, $T^{\prime}$ is the modified caption, and $A$ is the associated reference article.
For visual modification, we prompt GPT-4o to produce textual descriptions reflecting the intended visual manipulation. These descriptions are then used to generate images via the open-source FLUX.1 [dev] model Black Forest Labs (2024). The resulting sample is of the form $N_{\text{image}}=\{I^{\prime},T,A\}$, where $I^{\prime}$ is the modified image and $T$ is the original caption.
@@
For each of the 2,000 filtered news samples from Section 3.1, we generate misleading and non-misleading variants for each of the six creator intent categories outlined in Figure 2, resulting in a total of 12,000 instances. All samples are crafted to simulate the professional tone of real-world news reporting. Data examples are provided in Appendix A.4.
@@ 3.4 Human Verification of Data Quality
To assess the quality of news samples created in DeceptionDecoded, we conduct a human evaluation by randomly sampling 2% of the dataset, comprising 120 text and 120 image instances covering all six fine-grained intent categories. The goal is to determine whether human annotators perceive the misleading or non-misleading nature of each sample in alignment with the intent simulated by the news creators.
We recruit three graduate students as annotators. Each annotator is presented with 240 label-shuffled pairs of (original, DeceptionDecoded-generated) news pieces, where the generated version may be either misleading (M) or non-misleading (NM). Annotators are instructed to provide a binary label indicating whether the generated version sufficiently deviates from the original such that it could plausibly mislead an average reader. Additional details on the annotation protocol are provided in Appendix B.
We assess data quality using two metrics: (1) accuracy of labels in DeceptionDecoded (misleading/non-misleading) relative to the human-assigned label, based on aggregated human annotations, and (2) inter-annotator agreement, measured by Fleiss’ $\kappa$ Fleiss (1971). The results indicate high reliability and clarity, with 99.2% Acc. and $\kappa=0.877$ for text samples, and 89.2% Acc. with $\kappa=0.703$ for image samples.

### v3

# Seeing Through Deception: Uncovering Misleading Creator Intent in Multimodal News with Vision–Language Models

@@ Abstract
The impact of misinformation arises not only from factual inaccuracies but also from the misleading narratives that creators deliberately embed. Interpreting such creator intent is therefore essential for multimodal misinformation detection (MMD) and effective information governance. To this end, we introduce DeceptionDecoded, a large-scale benchmark of 12,000 image–caption pairs grounded in trustworthy reference articles, created using an intent-guided simulation framework that models both the desired influence and the execution plan of news creators. The dataset captures both misleading and non-misleading cases, spanning manipulations across visual and textual modalities, and supports three intent-centric tasks: (1) misleading intent detection, (2) misleading source attribution, and (3) creator desire inference. We evaluate 14 state-of-the-art vision–language models (VLMs) and find that they struggle with intent reasoning, often relying on shallow cues such as surface-level alignment, stylistic polish, or heuristic authenticity signals. These results highlight the limitations of current VLMs and position DeceptionDecoded as a foundation for developing intent-aware models that go beyond shallow cues in MMD.
Content Warning: this paper contains potentially harmful text and images.
@@ 1 Introduction
Multimodal misinformation, which combines persuasive text with compelling visuals, poses significant threats to public understanding and can lead to serious societal harm (Zhou & Zafarani, 2020; Alam et al., 2022; Do Nascimento et al., 2022). Research in multimodal misinformation detection (MMD) has primarily focused on cross-modal misalignment, typically in two forms: (1) out-of-context (OOC) misinformation (Yuan et al., 2023; Qi et al., 2024), where images and captions from unrelated events or time periods are falsely paired; and (2) multimodal media manipulation (Shao et al., 2023; Liu et al., 2025), involving subtle changes to visuals or captions that alter interpretation. Yet, current benchmarks rely on heuristic strategies such as CLIP-based mismatching (Luo et al., 2021) or sentiment substitutions (Sudhakar et al., 2019; Shao et al., 2023), which fail to capture the complexity of real-world multimodal misinformation.
@@
Beyond surface-level cross-modal misalignment, the key to effective real-world misinformation governance lies in detecting and understanding misleading creator intent (Appelman et al., 2022; Jaidka et al., 2025). Many misinformation campaigns are deliberately crafted to advance specific agendas, often without the audience’s awareness, yet still manage to substantially influence public opinion (Ecker et al., 2022; Broda & Strömbäck, 2024). As illustrated in Figure 1, even slight caption manipulations can significantly distort a trustworthy news context (e.g., by misattributing iceberg collapse to covert military actions) which ultimately erodes institutional trust. Although preliminary efforts have explored intent interpretation (Da et al., 2021; Gabriel et al., 2022; Wang et al., 2025), they are often limited to unimodal settings and fail to capture the synergy between visual and textual cues. Moreover, these approaches rely on reader inference of creator intent, which are prone to error due to the deceptive framing of such content. These limitations highlight a pressing need for transparent modeling of ground-truth creator intents in multimodal content, along with systematic evaluation frameworks for detecting misleading communicative objectives.
Figure 1: Why creator intent detection matters in multimodal news. Creators can distort trustworthy news contexts by crafting intent-loaded misinformation that appears semantically aligned (e.g., portraying broken icebergs) yet deliberately conveys false narratives (e.g., attributing them to secret underwater nuclear tests).
@@
To address this gap, we introduce an intent-guided framework for simulating intent-aware multimodal news reporting. In this framework, intent is defined as a combination of desired influence and execution plan, two components grounded in communication strategy theory (Hallahan et al., 2007; Paul, 2011). Leveraging this formulation, we construct DeceptionDecoded, a large-scale benchmark of 12,000 image–caption–article triplets $(I,T,A)$, where $I$ is the news image, $T$ is the accompanying caption, and $A$ is a trustworthy reference article (i.e., “trustworthy news context”) drawn from the VisualNews dataset of verified reporting (Liu et al., 2021). By anchoring manipulations to verifiable news contexts, this design enables explicit labeling of misleading versus non-misleading reporting, while also ensuring coverage of ten high-impact domains selected using journalistic and public-interest criteria. The benchmark further models realistic malicious influence scenarios such as political polarization and public health misinformation (Section 3.1), all while preserving a professional news reporting tone. Data quality and realism are validated through systematic human evaluation (Section 3.4), and the benchmark supports three intent-centric tasks central to effective MMD: (1) Misleading Intent Detection, which determines whether a multimodal news piece was deliberately crafted to mislead readers; (2) Misleading Source Attribution, which identifies whether the misleading signal originates from the image or the text; and (3) Creator Desire Inference, which infers the targeted area of societal impact (e.g., political polarization or public health disruption).
Using DeceptionDecoded, we benchmark 14 representative VLMs. Despite recent progress in multimodal reasoning (Hurst et al., 2024; Anthropic, 2025b), we find that state-of-the-art VLMs still struggle to recognize misleading creator intent in multimodal news. For instance, GPT-4o-mini achieves only 18.8% accuracy when intent is subtly encoded through caption manipulation (OpenAI, 2024). Models often rely on superficial cues such as image–text alignment, stylistic polish, or heuristic authenticity signals, leaving them vulnerable to simple adversarial manipulations including stylistic reframing (Chen & Shu, 2024; Wu et al., 2024) and persuasive prompting attacks (Zeng et al., 2024). Our analysis further reveals two broader insights. First, fine-tuning on DeceptionDecoded transfers improvements to external MMD benchmarks, demonstrating general utility. Second, advances in image generation improve visual fidelity, making manipulations more vivid, but VLMs still struggle with intent reasoning even under clearer visuals. At the same time, the scalability and realism of ever-evolving generators heighten the risk of persuasive misinformation, positioning DeceptionDecoded as both a principled benchmark and a stress test for developing intent-aware detectors that move beyond surface-level consistency.
@@
Benchmark Task Setup Content Modality Creator Intent Trustworthy Context
@@ 2 Related Work
2.1 Multimodal Misinformation Detection (MMD)
With the increasing prevalence of persuasive visual–textual misinformation on online platforms, research has expanded from detecting purely textual misinformation (Rashkin et al., 2017; Chen & Shu, 2024; Wu et al., 2024) to tackling multimodal misinformation detection (MMD). A growing body of work emphasizes misinformation arising from subtle cross-modal misalignments, such as out-of-context (OOC) misinformation (Qi et al., 2024) and multimedia manipulation (Shao et al., 2023; Liu et al., 2025). Recent advances leverage VLMs like LLaVA (Liu et al., 2023) and GPT-4o (Hurst et al., 2024) within retrieval-augmented frameworks, where the model reasons over multimodal inputs using retrieved trustworthy evidence (Khaliq et al., 2024; Xuan et al., 2024; Zhou et al., 2024; Braun et al., 2024). While effective in grounding responses, these approaches largely overlook the role of news creation intent, which has been identified as a key driver of misinformation (Sharma et al., 2019). Our work builds upon this evidence-based paradigm by evaluating VLMs on their ability to detect misleading communicative intent in multimodal news. We show that even the most advanced models still fall short in this setting, highlighting the need for intent-aware frameworks in MMD.
@@ 2.2 Misinformation Benchmarks
A variety of benchmarks have been developed to support misinformation-related research (Hanselowski et al., 2019; Yao et al., 2023). Representative datasets such as PolitiFact and GossipCop (Shu et al., 2020) collect naturally occurring misinformation from social media, while others approach multimodal misinformation detection (MMD) through multimedia manipulation; for example, NewsCLIPpings (Luo et al., 2021) and DGM4 (Shao et al., 2023) introduce cross-modal mismatches by substituting images or captions based on visual or textual similarity. Mixed-source datasets such as Fakeddit (Nakamura et al., 2020) and MMFakeBench (Liu et al., 2025) combine naturally occurring content with synthetically manipulated samples. Preliminary efforts such as EMU (Da et al., 2021), NewsINT (Wang et al., 2025), and MRF (Gabriel et al., 2022) have begun to consider news creator intent, but these remain limited to unimodal settings and typically define intent from readers’ subjective perspectives, which may diverge from the actual communicative purpose of the creator. In contrast, we introduce DeceptionDecoded, a benchmark of creator-produced multimodal misinformation where intent is explicitly defined during content generation. A comparison of DeceptionDecoded with prior datasets is shown in Table 1. Beyond this, DeceptionDecoded is guided by established communication theories, grounded in trustworthy news contexts, and validated through human evaluation, providing both a principled framework for modeling creator intent and a resource with broader utility for advancing general intent-aware MMD.
Figure 2: DeceptionDecoded: Overview of multimodal news curation guided by diverse simulated creator intents, covering both misleading and non-misleading cases.
@@ 3 DeceptionDecoded
To systematically investigate the role of misleading creator intent in multimodal news, it is necessary to construct a large-scale benchmark that (1) explicitly defines diverse communicative intents across multiple domains and (2) uses these intents to guide multimodal content generation. Achieving this with real-world news data is difficult, since published articles already reflect finalized narratives and annotators can only infer intent retrospectively from a reader’s perspective. To address this challenge, we introduce DeceptionDecoded (Figure 2), a benchmark of 12,000 multimodal news instances created through an intent-guided synthesis pipeline. Each instance is anchored in a trustworthy news context drawn from verified articles and paired with a predefined creator intent configuration, which enables controlled manipulation while maintaining content realism. The quality and plausibility of the generated data are further validated through rigorous human evaluation (see Section 3.4).
@@ 3.1 Source News Collection
We draw source data from VisualNews (Liu et al., 2021), a large-scale repository of trustworthy multimodal news, to ensure factual grounding. From this corpus, we select ten topic categories that provide broad societal relevance and sufficient sample richness, spanning domains such as politics, disasters, and public health. The full list of topics is provided in Appendix D.1.
To support meaningful evaluation of misinformation detection, we apply rigorous filtering criteria informed by both misinformation governance and professional journalism. From a governance perspective, we prioritize content that serves the public interest (Kruger et al., 2024), focusing on events with high societal relevance and potential for broad impact. Consistent with journalistic standards, we require that selected content demonstrates professionalism, neutrality, and clarity (Maras, 2013; Farley et al., 2014), ensuring that language remains objective, unbiased, and unambiguous. To further mitigate harm, we exclude instances referencing specific individuals or identifiable public figures (see Appendix D.2).
This process yields 2,000 high-quality news samples, evenly distributed across the ten categories. Each instance is represented as $N=\{I,T,A\}$, where $I$ is the news image, $T$ is the accompanying caption, and $A$ is the corresponding trusted reference article (i.e., “trustworthy news context”).
@@ 3.2 Creator Intent Establishment
Given the context provided by each original news instance $N=\{I,T,A\}$, we simulate both malicious and trustworthy content creators using GPT-4o (Hurst et al., 2024), leveraging its demonstrated capacity for social simulation (Anthis et al., 2025).
Inspired by strategic communication theories (Hallahan et al., 2007; Paul, 2011) and prior work on textual intent inference (Wang et al., 2025), we conceptualize creator intent ($C_{\text{int}}$) along two core dimensions: (1) desired influence and (2) execution plan. The desired influence specifies the societal sector(s) the creator seeks to affect (e.g., “public health and safety”), selected from a predefined taxonomy (up to three per instance; see Appendix D.3). The execution plan is expressed as open-ended text describing how the creator aims to achieve the intended influence.
@@ 3.3 Intent-Guided Multimodal News Creation
Guided by $C_{\text{int}}$, we generate both misleading and non-misleading variants of each news piece obtained in Section 3.1 by modifying either the image or the text. Misleading variants are categorized into two levels: (1) subtle, involving minor distortions to framing or background details that nudge reader interpretation, and (2) significant, involving major alterations that substantially change the perceived message. For non-misleading cases, the creator is prompted to faithfully paraphrase the caption or reconstruct the image while retaining consistency with the trustworthy news context $A$.
For textual modification, we use GPT-4o (Hurst et al., 2024) to generate captions aligned with the specified intent, as defined in Section 3.2, while keeping the original image unchanged. This produces samples of the form $N_{\text{text}}=\{I,T^{\prime},A\}$, where $I$ is the original image, $T^{\prime}$ is the modified caption, and $A$ is the corresponding trustworthy news context.
For visual modification, GPT-4o is prompted to generate textual descriptions of the intended manipulation, which are then used to synthesize images with the open-source FLUX.1 [dev] model (Black Forest Labs, 2024). The resulting sample takes the form $N_{\text{image}}=\{I^{\prime},T,A\}$, where $I^{\prime}$ is the modified image, $T$ is the original caption, and $A$ is the trustworthy news context.
@@
For each of the 2,000 filtered news samples from Section 3.1, we generate both misleading and non-misleading variants for each of the six creator intent categories outlined in Figure 2, yielding a total of 12,000 instances. All samples are crafted to reflect the professional tone of real-world news reporting. Illustrative data examples are provided in Appendix D.4.
@@ 3.4 Human Verification of Data Quality
To assess the quality of news samples created in DeceptionDecoded through intent-guided creation, we conduct a human evaluation by randomly sampling 2% of the dataset, comprising 120 text-modified and 120 image-modified instances that span all six fine-grained intent categories. The primary goal is to verify whether human annotators perceive the misleading or non-misleading nature of each sample in alignment with the intent simulated by the news creators.
We recruit three graduate students as annotators. Each annotator is presented with 240 label-shuffled pairs of (original, DeceptionDecoded-generated) news pieces, where the generated version may be either misleading (M) or non-misleading (NM). Annotators are instructed to provide a binary label indicating whether the generated version deviates from the original in a way that could plausibly mislead an average reader. Details of the annotation protocol are provided in Appendix E.
We evaluate data quality using two metrics: (1) accuracy, computed as agreement between the dataset labels and aggregated human annotations, and (2) inter-annotator agreement, measured by Fleiss’ $\kappa$ (Fleiss, 1971). Results show high reliability: 99.2% accuracy with $\kappa=0.877$ for text samples and 89.2% accuracy with $\kappa=0.703$ for image samples.
Beyond verifying the correctness of these binary labels, we also assess data realism and intent alignment. As described in Appendix E.2, annotators judged whether each synthesized piece could plausibly appear in today’s news or social media. Accuracy exceeded 98% for text and 91% for images in standalone plausibility (judging each piece in isolation), and remained above 87% in pairwise plausibility (judging original $\rightarrow$ manipulated pairs), confirming the contextual realism of the data. Furthermore, Appendix E.3 reports evaluations of whether generated manipulations remain faithful to their simulated desired influence and execution plan. Accuracy exceeded 92% for text and 86% for images across both dimensions, with substantial inter-annotator agreement.
Together, these evaluations demonstrate that DeceptionDecoded produces synthetic news samples that are not only correctly labeled as misleading or non-misleading, but also realistic in presentation and meaningfully aligned with the intended creator objectives.